OpenStudy (anonymous):

Write the quadratic equation in general form and then choose the value of "b." (2x - 1)(x + 5) = 0

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

[2x ^{2}+10x-x-5= 2x ^{2}+9x-5] the quadratic equation is 2x^2+9x-5 so b would be 9 not 11 this is because you have 5 times 2x= 10x and then -1 times x which is -x. when you take 10x+-x= 9x which makes b 9

hero (hero):

(2x - 1)(x + 5) = x(2x - 1) + 5(2x - 1) = 2x^2 - x + 10x - 5 = 2x^2 + 9x - 5 = 0